OpenStudy (anonymous):

Tina would like to melt down her existing gold jewelry to make a 20 ounce gold necklace that is 90% pure gold. After melting her jewelry, she is left with two batches. One batch is 75% pure. The other batch is 96% pure. How much of the 96% pure melted gold should Tina mix with the 75% pure melted gold to obtain 20 ounces of 90% pure gold? A. 5.71 ounces B. 7.00 ounces C. 14.29 ounces D. 13.00 ounces

OpenStudy (anonymous):

can i get help from someone pleaese

OpenStudy (mathstudent55):

Let x = amount of 75% pure gold Let y = amount of 96% pure gold The total amount of gold she will end up with is 20 ounces, so x + y = 20

OpenStudy (mathstudent55):

Now let's look at how much actual 100% pure gold she will end up with.

OpenStudy (anonymous):

do i divde 96 by all

OpenStudy (anonymous):

i dont get it

OpenStudy (mathstudent55):

x ounces of 75% pure gold contain 0.75x ounces of 100% gold. y ounces of 96% pure gold contain 0.96y ounces of 100% gold. The final result is 20 ounces of 90% gold, that means 0.90(20) = 18 ounces of 100% pure gold. This gives us our second equation.

OpenStudy (anonymous):

how did you get the 100%

OpenStudy (mathstudent55):

These are the two equations. x + y = 20 0.75x + 0.96y = 18

OpenStudy (mathstudent55):

Pure gold is 100% gold. When you say that gold is 75% pure, that means that by weight, 75% of it is gold, and 25% of it is not gold.

OpenStudy (anonymous):

oh ok so do i substitute the equations

OpenStudy (anonymous):

so half is not gold ok i get it a little bit so what do we do next

OpenStudy (mathstudent55):

Since we used the variable x to mean the amount of 75% gold, the question is how much pure gold is there in x ounces of 75% gold? The answer is 0.75x. Then we do the same for the 96% gold. 0.96y is the amount of pure gold in y ounces of 96% gold. Finally we do the same for the end result. The end result after mixing is 20 ounces of 90% gold. Here we actually know that the weight is 20 ounces, so the amount of pure gold in it is 90% of 20 oz or 0.9 * 20 = 18 ounces. Then we add the amounts of pure gold, and it must equal 18: 0.75x + 0.96y = 18 That's how we get the second equation.

OpenStudy (anonymous):

i got the wrong answer what did u get i got 14.29 ounces

OpenStudy (mathstudent55):

Ok, back to solving the problem. We have this system of equations. x + y = 20 0.75x + 0.96y = 18 We can use substitution. We can solve the first equation for one variable, let's say x, then substitute into the second equation.

OpenStudy (mathstudent55):

Which variable is equal to 14.29?

OpenStudy (anonymous):

i got it right yay thank you so much

OpenStudy (mathstudent55):

Great. You're welcome.

OpenStudy (anonymous):

did i get it right im not sure

OpenStudy (mathstudent55):

You didn't answer my question. Which variable did you get 14.29 ounces for? x or y?

OpenStudy (anonymous):

oh no i thought i got the answer right im sorry

OpenStudy (anonymous):

thank you so much

OpenStudy (mathstudent55):

x = 20 - y 0.75(20 - y) + 0.96y = 18 15 - 0.75y + 0.96y = 18 15 + 0.21y = 18 0.21y = 3 y = 14.29 Now substitute y = 14.29 in x + y = 20 to find x. x + 14.29 = 20 x = 5.71 Answer: 5.71 oz of 75% gold and 14.29 oz of 96% gold.

OpenStudy (mathstudent55):

You're welcome.